If the permutation of the word WHITE is selected at random, how
many of the permutations
i. Begins with a consonant?
ii. Ends with a vowel?
iii. Has a consonant and vowels alternating?

Answer by VFBundy(438)

i. Begins with a consonant?

3*4*3*2*1 = 72

ii. Ends with a vowel?

4*3*2*1*2 = 48

iii. Has a consonant and vowels alternating?

3*2*2*1*1 = 12